OpenStudy (anonymous):

What is the value of b2 - 4ac for the following equation? x(x + 8) = 9

OpenStudy (anonymous):

x(x+8)=9 \[x^{2}+8x-9=0\] a=1 b=8 c=-9 b^2=64 4*a*c=4*1*(-9) =-36 b^2-4ac=64-(-36) =64+36 =100